OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

Remember the slope of a line is always going to be equal. You're trying to find a point \((x, 0)\). So \[ \frac{-3-3}{1-(-5)}=\frac{-3-0}{1-x} \]

OpenStudy (anonymous):

So simplify it and then cross multiply to solve for \(x\).

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

Let's see. It simplifies to \[ \frac{-6}{6}=\frac{-3}{1-x}\implies 3=1-x\implies x=-2 \]

OpenStudy (anonymous):

It looks like you got it right.